### v2

# Geometry of horospheres in Kobayashi hyperbolic domains

Abstract.
For a Kobayashi hyperbolic domain, Abate introduced the notion of small and big horospheres of a given radius at a boundary point with a pole. In this article, we investigate which domains have the property that closed big horospheres and closed small horospheres centered at a given point and of a given radius intersect the boundary only at that point? We prove that any model-Gromov-hyperbolic domain have this property. To provide examples of non-Gromov-hyperbolic domains, we show that unbounded locally model-Gromov-hyperbolic domains and bounded, Dini-smooth, locally convex domains — that are locally visible — also have this property. Finally, using the geometry of the horospheres, we present a result about the homeomorphic extension of biholomorphisms and give an application of it.
Key words and phrases:
Kobayashi distance, geodesics, horospheres, Gromov product, Gromov hyperbolicity, visibility, Dini-smooth boundary
2020 Mathematics Subject Classification:
Primary: 32F45, 53C23; Secondary: 32Q45, 32A40, 53C22.
1. Introduction and statement of the results
Let $\Omega\subset\mathbb{C}^{d}$ be a Kobayashi hyperbolic domain equipped with the Kobayashi distance $\mathsf{k}_{\Omega}$. Given a point $x\in\partial\Omega$, $o\in\Omega$ and $R>0$, Abate [1] defined the following objects:
$H^{s}_{o}(x,R)$ $:=\Big\{z\in\Omega\,:\,\limsup_{w\to x}\big(\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(o,w)\big)<(1/2)\log R\Big\}$
$H^{b}_{o}(x,R)$ $:=\Big\{z\in\Omega\,:\,\liminf_{w\to x}\big(\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(o,w)\big)<(1/2)\log R\Big\}.$
The sets $H^{s}_{o}(x,R),H^{b}_{o}(x,R)$ are called the small horosphere and the big horosphere with center $x$, pole $o$ and radius $R$ respectively. Some of their basic properties are stated in Section 2. Note that although the objects $H^{s}_{o}(x,R)$, $H^{b}_{o}(x,R)$ can be defined for any Kobayashi hyperbolic domain, they could be empty set. For $H^{b}_{o}(x,R)$ completeness of $\mathsf{k}_{\Omega}$ is a natural condition under which they are nonempty for all choices of parameters; see Lemma 3.2 and Lemma 3.1 in Section 3. On the other hand, $H^{s}_{o}(x,R)$ could be empty, for some choice of parameters, for a domain $\Omega$ with $\mathsf{k}_{\Omega}$ complete; see e.g. Example 5.3.
In the case $\Omega=\mathbb{D}$, the open unit disc in the complex plane centered at 0, using the formula for $\mathsf{k}_{\mathbb{D}}$, a straightforward computation shows that when $o=0$, $x\in\partial\mathbb{D}$, for each $z\in\mathbb{D}$, one has
$\lim_{w\to x}\big(\mathsf{k}_{\mathbb{D}}(z,w)-\mathsf{k}_{\mathbb{D}}(o,w)\big)=\frac{1}{2}\log\Bigg(\frac{|x-z|^{2}}{1-|z|^{2}}\Bigg),$
and therefore
$H^{s}_{o}(x,R)=H^{b}_{o}(x,R)=H(x,R):=\Bigg\{z\in\mathbb{D}\,:\,\frac{|x-z|^{2}}{1-|z|^{2}}<R\Bigg\}.$
Geometrically, $H(x,R)$ is the disc with center x/(1+R)x/{(1+R)} and radius $R/(R+1)$, internally tangent to $\partial\mathbb{D}$ at $x$. In particular, closures of both the small and big horospheres of arbitrary radius with center $x$ and pole at 0 intersect the boundary of the unit disc only at the point $x$. An analogous result holds true for the Euclidean unit ball $\mathbb{B}^{n}$ of $\mathbb{C}^{n}$, $n\geq 2$. On the other hand, for the $n$-dimensional polydisc ${\mathbb{D}}^{n}$, $n\geq 2$, things are quite different. In this case, taking $o=0\in{\mathbb{D}}^{n}$, Abate showed that $\overline{H^{s}_{0}(x,R)}\cap\partial{\mathbb{D}}^{n}=\{x\}$ if and only if $x$ is on the Shilov boundary $(\partial\mathbb{D})^{n}$ of ${\mathbb{D}}^{n}$. On the other hand, $\overline{H^{b}_{0}(x,R)}\cap\partial{\mathbb{D}}^{n}$ always contains $\{x\}$ properly (see Proposition 2.4.12 and Corollary 2.4.13 in [2]).
Motivated from the above observations, we ask the following problem:
    Under what conditions on a domain $\Omega\subset\mathbb{C}^{d}$, it is true that for any $x\in\partial\Omega$, $o\in\Omega$ and $R>0$
    $\overline{H^{s}_{o}(x,R)}\cap\partial{\Omega}=\overline{H^{b}_{o}(x,R)}\cap\partial{\Omega}=\{x\}?$ (1.1)
Geometrically, condition 1.1 above implies that both $H^{s}_{o}(x,R)$ and $H^{b}_{o}(x,R)$ should be nonempty and their diameter tends to zero as $R$ decreases to zero. If we restrict our attention to bounded simply-connected planar domains, to the best of our knowledge, we are not aware of any result that characterizes such domains having property 1.1 above. We present such a characterization.
Theorem 1.1.
Let $\Omega\subset\mathbb{C}$ be a bounded simply-connected domain. Let $\psi:\mathbb{D}\longrightarrow\Omega$ be a Riemann map and let $o=\psi(0)$. Then the following are equivalent.
  • For any $x\in\partial\Omega$ and for any $R>0$, $\overline{H^{s}_{o}(x,R)}\cap\partial\Omega=\overline{H^{b}_{o}(x,R)}\cap\partial\Omega=\{x\}$.
  • The map $\psi:\mathbb{D}\longrightarrow\Omega$ extends as a homeomorphism from $\overline{\mathbb{D}}$ onto $\overline{\Omega}$, i.e., $\Omega$ is a Jordan domain.
An upshot of the above theorem is that for bounded planar simply-connected domains condition 1.1 above implies that the small horospheres and big horospheres with the same parameters coincide. Theorem 1.1 is a consequence of one of the main results of this paper, namely Theorem 1.9, regarding the homeomorphic extension of a biholomorphism between bounded domains. This later result gives another reason of studying domains having condition 1.1.
In higher dimensions, Abate first proved that for a bounded strongly pseudoconvex domain property (1.1) holds true; see [2, Theorem 2.4.14]. In fact, using Lempert’s theory of complex geodesics [18], Abate also showed that if the boundary of a bounded strongly convex domain is at least $\mathcal{C}^{3}$, then, for each $z\in\Omega$, the limit
$\lim_{w\to x}\big(\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(o,w)\big)$ (1.2)
exists and thus as a consequence big and small horospheres coincide (see [3]).
To understand the existence of limit in (1.2), Arosio et al. in [4] considered the horofunction compactification $\overline{\Omega}^{H}$ and the horofunction boundary $\partial_{H}\Omega$. These notions were introduced by Gromov in 1981 [17]. Later, Marc Rieffel rediscovered this compactification as the maximal ideal space of a unital cummutative $C^{*}$-algebra. He refers to it as the metric compactification and the corresponding boundary as the metric boundary; see [22, Section 4]. From this later point of view, for a bounded domain $\Omega$, the existence of the limit in (1.2) is equivalent to the existence of a continuous map from $\overline{\Omega}$ to $\overline{\Omega}^{H}$ extending the identity map on $\Omega$. Working in this setting, Arosio et al. proved that if $\Omega$ is either
  • a bounded strongly pseudoconvex domain with $\mathcal{C}^{2}$-boundary in $\mathbb{C}^{d}$,
  • a bounded convex domain such that $s_{\Omega}(z)\to 1$ as $z\to\partial\Omega$,
  • a bounded convex domain in $\mathbb{C}^{d}$ with $\mathcal{C}^{\infty}$-boundary, of finite type in the sense of D’Angelo,
then the limit in (1.2) exists. Here, $s_{\Omega}(z)$ denotes the squeezing function of the domain; see [16] for the definition and basic properties.
The above three types of domains are model-Gromov-hyperbolic domains, i.e., $(\Omega,\mathsf{k}_{\Omega})$ is Gromov hyperbolic and the identity map on $\Omega$ extends as a homeomorphism from the Gromov compactification $\overline{\Omega}^{G}$ to $\overline{\Omega}$; this follows from the works of Balogh–Bonk [5], Zimmer [26, 25], together with the results in [9]. Arosio et al. observed that in the above three cases the distance space $(\Omega,\mathsf{k}_{\Omega})$ has an additional property, namely, the approaching geodesic property. This property gives a sufficient condition under which the identity map on $\Omega$ extends as a homeomorphism from $\overline{\Omega}^{G}$ to $\overline{\Omega}^{H}$, i.e., any two of the three compactifications $\overline{\Omega}$, $\overline{\Omega}^{G}$, $\overline{\Omega}^{H}$ admits homeomorphisms extending the identity map.
Recently, Bharali–Zimmer [7, Theorem 1.11] showed that a model-Gromov-hyperbolic domain is Cauchy complete and satisfies visibility property with respect to the geodesics of the Kobayashi distance, see 2.2 for the definition of this property. In this paper, we show that
Proposition 1.2.
Let $\Omega$ be a complete Kobayashi hyperbolic domain that satisfies the visibility property with respect to the geodesics of the Kobayashi distance. Then for any $o\in\Omega$, $x\in\partial\Omega$ and $R>0$ we have
$\overline{H_{o}^{b}(x,R)}\cap\partial{\Omega}=\{x\}.$
As an upshot of the above result and the existence of limit in (1.2), for the three classes of domains considered by Arosio et al., property (1.1) in the problem ()(*) holds true. However, whether (1.1) holds true for a general model-Gromov-hyperbolic domain is not known. In this article, we answer this question affirmatively. Before we state this result, a notation: since we would also like to consider unbounded domains, the natural formulation of model-Gromov-hyperbolic domains is via the end-point compactification $\overline{\Omega}^{End}$. For bounded domains $\overline{\Omega}^{End}=\overline{\Omega}$; see [14] for more details about end-point compactification.
Theorem 1.3.
Let $\Omega$ be a Kobayashi hyperbolic domain and suppose $(\Omega,\mathsf{k}_{\Omega})$ is Gromov hyperbolic metric space. If the identity map ${\sf id}_{\Omega}$ extends to a homeomorphism from $\overline{\Omega}^{G}$ onto $\overline{\Omega}^{End}$ then for any $x\in\partial{\Omega}$, $o\in\Omega$ and $R>0$, we have:
$\overline{H^{s}_{o}(x,R)}\cap\partial\Omega=\overline{H^{b}_{o}(x,R)}\cap\partial\Omega=\{x\}.$
Owing to the aforementioned result of Bharali–Zimmer and Proposition 1.2, we know that $\overline{H^{b}_{o}(x,R)}\cap\partial\Omega=\{x\}$ for every $x\in\partial\Omega$, for a model-Gromov-hyperbolic domain. Gromov hyperbolicity plays a crucial role in showing that $x\in\overline{H^{s}_{o}(x,R)}$. More specifically, it is the property of uniform separation of geodesic rays starting at a fixed point in the domain and landing at the same point in the Gromov boundary that is at the heart of our proof that $\overline{H^{s}_{o}(x,R)}\cap\partial\Omega=\{x\}$. There is no complete characterization of a model-Gromov-hyperbolic domain. Except the three considered by Arosio et al., a few others are also listed in [11].
Remark 1.4.
We also present an alternative proof — suggested by the referee — that does not rely on visibility. This second proof draws upon the horofunction compactification, the correspondence between the horofunction and Gromov boundaries together with the uniform boundedness of horofunctions corresponding to the same Gromov boundary point. The two proofs emphasize different geometric aspects, namely — visibility versus the horofunction compactification.
A natural question arises whether there are non-Gromov-hyperbolic domains having property (1.1) in the problem ()(*). We provide two results in this direction. Our first result deals with domains that are locally model-Gromov-hyperbolic domain; see Section 4 for the definition.
Theorem 1.5.
Let $\Omega\subset\mathbb{C}^{d}$ be an complete Kobayashi hyperbolic domain that is hyperbolically embedded in $\mathbb{C}^{d}$. Suppose $\Omega$ is locally model-Gromov-hyperbolic domain. Then for any $x\in\partial\Omega$, $o\in\Omega$ and $R>0$, we have
$\overline{H^{s}_{o}(x,R)}\cap\partial{\Omega}=\overline{H^{b}_{o}(x,R)}\cap\partial{\Omega}=\{x\}.$
We refer the reader to Subsection 2.2 for the definition of a domain being hyperbolically embedded in $\mathbb{C}^{d}$. A recent result of Bracci et al. [11] says that any bounded local model-Gromov-hyperbolic domain is itself a model-Gromov-hyperbolic domain. In Section 6, we shall provide a class of planar domains, introduced in [14], that are locally model-Gromov-hyperbolic, however, in general, they are not Gromov hyperbolic domains.
The other class of domains that satisfy condition (1.1) and provides examples of non-Gromov-hyperbolic bounded domains are modelled locally on bounded convex domains with Dini-smooth boundary, see [21] for the definition. To state the result we need a definition. Let $\Omega$ be a domain in $\mathbb{C}^{d}$. We say $\Omega$ is locally convex if for any point $x\in\partial\Omega$, there exist $\mathbb{C}^{d}$-neighbourhoods $U_{x}$, $V_{x}$ of $x$ such that $V_{x}\Subset U_{x}$, and an injective holomorphic map $\Phi_{x}:U_{x}\longrightarrow\mathbb{C}^{d}$ such that $\Phi_{x}(\Omega\cap V_{x})$ is an open convex set.
Theorem 1.6.
Let $\Omega$ be a complete Kobayashi hyperbolic bounded domain with Dini-smooth boundary. Suppose $\Omega$ is locally convex. Moreover, suppose for any point $x\in\partial\Omega$, the domain $\Phi_{x}(V_{x}\cap\Omega)$ has visibility property with respect to the geodesics, where $(\Phi_{x},U_{x},V_{x})$ be as in the definition of local convexity. Then for any $x\in\Omega$, $o\in\Omega$ and $R>0$, we have
$\overline{H^{s}_{o}(x,R)}\cap\partial{\Omega}=\overline{H^{b}_{o}(x,R)}\cap\partial{\Omega}=\{x\}.$
A bounded convex domain with smooth boundary having all the points in its boundary — except a nonempty finite subset — of finite D’Angelo type is a visibility domain, see [10, Theorem 1.1], [13, Corollary 1.4]. So the above theorem applies to it. Note that by a result of Zimmer [25], such a domain is not Gromov hyperbolic. Bharali–Zimmer [6] also constructed bounded convex domains — that are not Gromov hyperbolic — and has visibility property. The above theorem applies to such domains too. A recent result of Bracci–Nikolov–Thomas says that any complete Kobayashi hyperbolic bounded domain with Dini-smooth boundary such that all its boundary points are $\mathbb{C}$-strictly convex is a visibility domain; see [10, Theorem 1.3]. Theorem 1.6 applies to all such domains too.
The domains in Theorem 1.3, Theorem 1.5 and Theorem 1.6 are all visible domains. It is natural to ask if the visibility property itself implies the property (1.1) in problem ()(*). This is not true, e.g., the domain $\Omega:=\mathbb{D}\setminus\{[0,1)\}$ is a visibility domain but for any $x\in(0,1]$, there exists an $R_{0}$, depending on $x$, such that for any $R<R_{0}$, we have
$H^{s}_{o}(x,R)=\emptyset\ \ \forall R<R_{0}$
and for some $o\in\Omega$; see Example 5.3 for the details.
We now present our result regarding the continuous extension of biholomorphisms. To present our results, we define a class of ‘nice domains’. Given a complete Kobayashi hyperbolic domain $D$, let $\xi\in\partial{D}$ and $o\in D$ be such that for any $z\in D$ the limit
$\lim_{w\to\xi}\big(\mathsf{k}_{D}(z,w)-\mathsf{k}_{D}(o,w)\big)\ \ \text{exists}.$
Then, for all $R>0$, we define the horosphere centered at $\xi$ of radius $R$ and pole $o$ as
$H_{o}(\xi,R):=\big\{z\in D\,:\,\lim_{w\to\xi}\big(\mathsf{k}_{D}(z,w)-\mathsf{k}_{D}(o,w)\big)<(1/2)\log R\big\}.$
When the context is clear, we shall write $H_{o}(\xi,R)$ simply as $H(\xi,R)$.
Definition 1.7.
A complete Kobayashi hyperbolic domain $D$ is said to be metrically-regular if there exists a point $o\in D$ such that
  1. (1)
    for any $\xi\in\partial D$ and $z\in D$, the limit
    $\lim_{w\to\xi}\big(\mathsf{k}_{D}(z,w)-\mathsf{k}_{D}(o,w)\big)\ \ \text{exists and}$
  2. (2)
    given two distinct points $\xi_{1}\neq\xi_{2}\in\partial{D}$, there exists an $R>0$ such that
    $H(\xi_{1},R)\cap H(\xi_{2},R)=\emptyset.$
The Euclidean unit ball $\mathbb{B}^{d}\subset\mathbb{C}^{d}$, $d\geq 1$, is metrically-regular as one explicitly knows the formulae for the Kobayashi distance. On the other hand, $\mathbb{D}^{d}$, $d\geq 2$, is not metrically-regular. The three types of domains considered by Arosio et al. are metrically-regular.
Remark 1.8.
It is evident that when $\Omega$ is bounded, the metric regularity of $\Omega$ implies that the identity map on $\Omega$ extends as a homeomorphism from its horofunction compactification $\overline{\Omega}^{H}$ to its Euclidean closure $\overline{\Omega}$.
We are now ready to present our main result.
Theorem 1.9.
Let $D$, $\Omega$ be bounded domains that are complete with respect to the Kobayashi distance. Suppose $D$ is metrically-regular and $\Omega$ has the property that
  • $a)$
    for each $x\in\partial\Omega$ and for every $R>0$, $H^{s}_{p}(x,R)$ is nonempty with respect to some (and hence any) pole $p$, and
  • $b)$
    for any two distinct points $x,y\in\partial{\Omega}$ there is an $R>0$ such that $H^{b}_{p}(x,R)\neq H^{b}_{p}(y,R)$.
Then any biholomorphism $\Psi:D\longrightarrow\Omega$ extends as a homeomorphism from $\overline{D}$ onto $\overline{\Omega}$.
Remark 1.10.
Note that any bounded domain that satisfies condition (1.1) in problem ()(*) satisfies the condition $(a)$ and $(b)$ imposed on the domain $\Omega$ in the statement of the above theorem. In particular, taking $D=\mathbb{D}$ and $\Omega$ to be a bounded simply-connected domain in $\mathbb{C}$ that satisfies (1.1) we recover Theorem 1.1. Theorem 1.9 also implies that bounded simply-connected domains that are metrically-regular are the Jordan domains.
2. Preliminaries
2.1. Basic properties of horospheres.
We begin this section with a list of important properties of the horospheres due to Abate, see [Section 2.4.2][2]. In what follows, $B_{\mathsf{k}}(o,r)$ will denote the ball with respect to the Kobayashi distance centered at $o$ of radius $r>0$.
Result 2.1 (paraphrasing Lemma 2.4.10 in [2]).
Let $\Omega$ be a Kobayashi hyperbolic domain. Then for any $o,p\in\Omega$, $x\in\partial{\Omega}$, we have
  1. (1)
    for every $R>0$ we have $H^{s}_{o}(x,R)\subset H^{b}_{o}(x,R)$;
  2. (2)
    for every $0<R_{1}<R_{2}$, $H^{s}_{o}(x,R_{1})\subset H^{s}_{o}(x,R_{2})$ and $H^{b}_{o}(x,R_{1})\subset H^{b}_{o}(x,R_{2})$;
  3. (3)
    for every $R>1$, we have $B_{\mathsf{k}}(o,(1/2)\log R)\subset H^{s}_{o}(x,R)$;
  4. (4)
    for every $R<1$ we have $H^{b}_{o}(x,R)\cap B_{\mathsf{k}}(o,(-1/2)\log R)=\emptyset$;
  5. (5)
    $\bigcup_{R>0}H^{s}_{o}(x,R)=\bigcup_{R>0}H^{b}_{o}(x,R)=\Omega$ and $\bigcap_{R>0}H^{s}_{o}(x,R)=\bigcap_{R>0}H^{b}_{o}(x,R)=\emptyset$.
  6. (6)
    for any $R>0$, we have ${H_{p}^{b}(x,R)}\subset{H_{o}^{b}(x,LR)}$, ${H_{p}^{s}(x,R)}\subset{H_{o}^{s}(x,LR)}$ where
    $({1}/{2})\log L:=\limsup_{w\to x}\big(\mathsf{k}_{\Omega}(p,w)-\mathsf{k}_{\Omega}(o,w)\big).$
When there are more than one domains under consideration, we shall use $H^{s}_{o}(x,R,\Omega)$, $H^{b}_{o}(x,R,\Omega)$ to specify the domain in which horospheres are considered.
2.2. Results regarding visibility property of a domain
In this subsection, we put together certain results regarding visibility property of a Kobayashi hyperbolic domain.
Definition 2.2.
A complete Kobayashi hyperbolic domain $\Omega\subset\mathbb{C}^{d}$ is said to possess visibility property with respect to the geodesics of the Kobayashi distance $\mathsf{k}_{\Omega}$ if for any two distinct points $x$, $y\in\partial{\Omega}$, there exist disjoint $\mathbb{C}^{d}$-neighbourhoods $U_{x}$, $U_{y}$ of $x$ and $y$ respectively, and a compact subset $K\subset\Omega$ such that any geodesic that starts in $U_{x}$ and ends in $U_{y}$ intersects $K$.
In this paper, we shall mostly be concerned with this form of visibility which is designed for complete spaces. There are two other notions of visibility, namely weak visibility and visibility that are designed considering that, in general, it is difficult to determine whether $(\Omega,\mathsf{k}_{\Omega})$ is complete. The definition of weak visibility is analogous to Definition 2.2 except that the geodesics are replaced by $(1,\kappa)$-almost geodesics for all $\kappa>0$. For more information, we refer the reader to [6], [10], [7]. For complete hyperbolic domains, we have the following:
Result 2.3 ([10, Proposition 2.5], [13, Proposition 3.1]).
Let $\Omega$ be a complete Kobayashi hyperbolic domain in $\mathbb{C}^{d}$. Then the following are equivalent.
  • $\Omega$ has visibility property with respect to the geodesics.
  • $\Omega$ has weak visibility property.
  • For any $o\in\Omega$, and for any $x\neq y\in\partial\Omega$ we have
    $\limsup_{z\to x,\,w\to y}\langle z\,|\,w\rangle^{\Omega}_{o}<\infty.$
Here, $\langle z\,|\,w\rangle^{\Omega}_{o}$ denotes the Gromov product of $z$ and $w$ with respect to $o$ defined by
$\langle z\,|\,w\rangle^{\Omega}_{o}:=\frac{\mathsf{k}_{\Omega}(o,z)+\mathsf{k}_{\Omega}(o,w)-\mathsf{k}_{\Omega}(z,w)}{2}.$
The above result was established for bounded domains. However, it can also be proved for unbounded domains using a ‘localization trick’ provided by [14, Lemma 2.2] to prove [14, Theorem 1.3].
A Kobayashi hyperbolic domain $\Omega\subset\mathbb{C}^{d}$ is called hyperbolically embedded in $\mathbb{C}^{d}$ if for any two distinct points $x\neq y\in\partial\Omega$, we have
$\liminf_{z\to x,\,w\to y}\mathsf{k}_{\Omega}(z,w)>0.$
It is a fact that all planar domains are hyperbolically embedded in $\mathbb{C}$. A bounded domain $\Omega\subset\mathbb{C}^{d}$ is hyperbolically embedded in $\mathbb{C}^{d}$. We now state the next result.
Result 2.4 ([14, Theorem 1.6], [19, Theorem 1.6]).
Let $\Omega\subset\mathbb{C}^{d}$ be a Kobayashi hyperbolic domain that is hyperbolically embedded in $\mathbb{C}^{d}$. Suppose for any $x\in\partial\Omega$, there exists a $\mathbb{C}^{d}$-neighbourhood $U_{x}$ of $x$ such that $U_{x}\cap\Omega$ is connected and $(U_{x}\cap\Omega,\mathsf{k}_{U_{x}\cap\Omega})$ has weak visibility property. Then $(\Omega,\mathsf{k}_{\Omega})$ has weak visibility property.
The last result that we state is a localization result for the Kobayashi distance in weak visibility domains.
Result 2.5 (Corollary of Theorem 1.3 in [23]).
Let $\Omega$ be a Kobayashi hyperbolic domain that has weak visibility property. Suppose for any point $x\in\partial{\Omega}$, there exist $\mathbb{C}^{d}$-neighbourhoods $U_{x}$, $W_{x}$ of $x$ such that $W_{x}\Subset U_{x}$ and $\Omega\cap U_{x}$ is connected. Then there exists a $C>0$ such that
$\mathsf{k}_{U_{x}\cap\Omega}(z,w)\leq\mathsf{k}_{\Omega}(z,w)+C,\ \ \forall z,w\in W_{x}\cap\Omega.$
3. General results about horospheres
In this section, we present some general results about horospheres. At the end of this section, we shall present the proof of Proposition 1.2. We begin with an elementary lemma.
Lemma 3.1.
Let $\Omega$ be a hyperbolic domain and let $o\in\Omega$ and $x\in\partial{\Omega}$ be given. Suppose that
$M:=\limsup_{w\to x}\mathsf{k}_{\Omega}(o,w)<\infty.$
Then for any $0<R<e^{-2M}$, we have $H^{s}_{o}(x,R)\subset H^{b}_{o}(x,R)=\emptyset$.
Proof.
For a fixed $z\in\Omega$, we have $\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(o,w)\geq-\mathsf{k}_{\Omega}(o,w)$. Therefore
$\liminf_{w\to x}\big(\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(o,w)\big)\geq\liminf_{w\to x}\big(-\mathsf{k}_{\Omega}(o,w)\big)=-\limsup_{w\to x}\mathsf{k}_{\Omega}(o,w)=-M$
Now if $z\in H^{b}_{o}(x,R)$ for some $R>0$, then the above inequality and the definition of $H^{b}_{o}(x,R)$ implies
$M+{1}/{2}\log R>0\Longleftrightarrow R>e^{-2M}.$
This implies that for any $R<e^{-2M}$ the big horosphere $H^{b}_{o}(x,R)=\emptyset$. ∎
The above lemma suggests that to study the geometry of the horospheres with respect to a pole $o\in\Omega$, we should impose on $\Omega$ the following condition:
$\limsup_{w\to x}\mathsf{k}_{\Omega}(o,w)=\infty\ \ \text{$\forall x\in\partial{\Omega}$}.$
Note that every complete Kobayashi hyperbolic domain satisfies this property. Indeed for such domains we have the following:
Lemma 3.2.
Let $\Omega$ be a complete Kobayashi hyperbolic domain in $\mathbb{C}^{d}$. Then for any $o\in\Omega$, $x\in\partial{\Omega}$ and $R>0$, the set $H^{b}_{o}(x,R)$ is non-empty. Moreover, if $\Omega$ is bounded, then the set $\overline{H^{b}_{o}(x,R)}\cap\partial{\Omega}$ is also non-empty.
Proof.
Fix $o\in\Omega$, $x\in\partial{\Omega}$ and let $R>0$ be given. Choose a sequence $(w_{n})\subset\Omega$ such that $w_{n}\to x$ as $n\to\infty$ and, for a given $s>0$, consider the ball $B_{\mathsf{k}}(o,s)$. Since $(\Omega,\mathsf{k}_{\Omega})$ is a locally compact length metric space that is complete, by Hopf–Rinow Theorem, every bounded set is relatively compact. Therefore, without loss of generality, we can assume that $w_{n}\in\Omega\setminus B_{\mathsf{k}}(o,s)$. For any $n$, choose $b_{n}\in\partial B_{\mathsf{k}}(o,s)$ such that
$\mathsf{k}_{\Omega}(o,w_{n})=\mathsf{k}_{\Omega}(o,b_{n})+\mathsf{k}_{\Omega}(b_{n},w_{n}).$ (3.1)
By passing to subsequences, we may assume that $b_{n}\to b_{s}$ as $n\to\infty$ for some $b_{s}\in\partial B_{\mathsf{k}}(o,s)$. Therefore, given $\epsilon>0$ there exists $n_{0}\in\mathbb{N}$ such that
$\mathsf{k}_{\Omega}(b_{s},b_{n})<\epsilon\ \ \ \forall n\geq n_{0}.$
Claim. $b_{s}\in H^{b}_{o}(x,e^{-s})$.
Given $\epsilon>0$ the triangle inequality and the inequality above implies
$\begin{split}\mathsf{k}_{\Omega}(b_{s},w_{n})&\leq\mathsf{k}_{\Omega}(b_{s},b_{n})+\mathsf{k}_{\Omega}(b_{n},w_{n})\\ &<\epsilon+\mathsf{k}_{\Omega}(b_{n},w_{n})\hskip 8.53581pt\forall n\geq n_{0}.\end{split}$
The above inequality and (3.1) implies that, for all $n\geq n_{0}$, we have
$\mathsf{k}_{\Omega}(b_{s},w_{n})-\mathsf{k}_{\Omega}(o,w_{n})<-s+\epsilon.$
Therefore, for every $\epsilon>0$, we have
$\liminf_{w\to x}\big(\mathsf{k}_{\Omega}(b_{s},w)-\mathsf{k}_{\Omega}(o,w)\big)\leq\liminf_{n\to\infty}\big(\mathsf{k}_{\Omega}(b_{s},w_{n})-\mathsf{k}_{\Omega}(o,w_{n})\big)\leq-s+\epsilon.$
This implies that
$\liminf_{w\to x}\big(\mathsf{k}_{\Omega}(b_{s},w)-\mathsf{k}_{\Omega}(o,w)\big)<({1}/{2})\log{e}^{-s}.$
[…]
It is enough to show that $H^{b}_{o}(x,R)$ is non-empty for all $R<1$. Fix such an $R$, then for every $s\in(-\log{R},\infty)$, there exists a point $b_{s}$ such that $\mathsf{k}_{\Omega}(o,b_{s})=s$ and $b_{s}\in H^{b}_{o}(x,e^{-s})\subset H^{b}_{o}(x,R)$. This implies that $H^{b}_{o}(x,R)$ is nonempty. If $\Omega$ is bounded, since $\lim_{s\to\infty}\mathsf{k}_{\Omega}(o,b_{s})=\infty$, the set $\overline{H^{b}_{o}(x,R)}\cap\partial{\Omega}$ is nonempty too, for any $R<1$, and hence for any $R>0$. ∎
The above lemma does not give any information about the points in the set $\overline{H^{b}(x,R)}\cap\partial{\Omega}$. Our next lemma is a result in this direction.
Lemma 3.3.
Let $\Omega$ be a complete Kobayashi hyperbolic domain and let $\gamma:[0,\infty)\longrightarrow(\Omega,\mathsf{k}_{\Omega})$ be a geodesic ray. Suppose $x,y\in\partial{\Omega}$ be such that there exist sequences $(t_{n})$ and $(s_{n})$ such that $t_{n}\to\infty$, $s_{n}\to\infty$ and
$x=\lim_{n\to\infty}\gamma(t_{n}),\ \ \ y=\lim_{n\to\infty}\gamma(s_{n}).$
Then, for all $R>0$, $o\in\Omega$, $y\in\overline{H^{b}_{o}(x,R)}$; in particular, taking $s_{n}=t_{n}$, $x\in\overline{H^{b}_{o}(x,R)}$.
Proof.
Let $p=\gamma(0)$.
Claim. For all $R>0$, $y\in\overline{H_{p}^{b}(x,R)}$.
Suppose $y\notin\overline{H_{p}^{b}(x,R)}$ for some $R>0$. Then there exists an $r>0$ such that for any $z\in B(y,r)\cap\Omega$ we have
$\liminf_{w\to x}\big(\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(p,w)\big)\geq(1/2)\log R.$
Since $\gamma(s_{n})\to y$, there exists $N_{0}\in\mathbb{N}$ such that for all $n\geq N_{0}$, $\gamma(s_{n})\in B(y,r)$. Fix a $k\geq N_{0}$, then as $\gamma(t_{n})\to x$, we have
$\liminf_{n\to\infty}\big(\mathsf{k}_{\Omega}(\gamma(s_{k}),\gamma(t_{n}))-\mathsf{k}_{\Omega}(\gamma(0),\gamma(t_{n}))\big)$ $\geq\liminf_{w\to x}\big(\mathsf{k}_{\Omega}(\gamma(s_{k}),w)-\mathsf{k}_{\Omega}(\gamma(0),w)\big)$
$\geq(1/2)\log R.$
Since $\gamma$ is a geodesic ray, the above inequality implies that
$s_{k}+(1/2)\log R\leq 0\ \ \ \forall k\geq N_{0},$
[…]
By Lemma 2.1, part $(6)$, for all $S>0$, we have ${H_{p}^{b}(x,S)}\subset{H_{o}^{b}(x,LS)}$ where $L$ is given by the identity
$({1}/{2})\log L=\limsup_{w\to x}\big(\mathsf{k}_{\Omega}(p,w)-\mathsf{k}_{\Omega}(o,w)\big).$
Clearly if $L\leq 1$, then we are done, otherwise for a given $R$ we choose $S>0$ small enough such that $LS<R$, since $y\in\overline{H_{p}^{b}(x,S)}$, the inclusion above implies the desired conclusion. ∎
Given a geodesic ray $\gamma:[0,\infty)\longrightarrow(\Omega,\mathsf{k}_{\Omega})$, consider the cluster set of $\gamma$
$\mathcal{C}(\gamma):=\big\{x\in\partial{\Omega}\,:\,\text{$\exists$ a sequence $(t_{n})$, $t_{n}\to\infty$, such that $\gamma(t_{n})\to x$}\big\}.$
Lemma 3.3 says that for any $x\in\mathcal{C}(\gamma)$, $\mathcal{C}(\gamma)\subset\overline{H^{b}_{o}(x,R)}$ for all $R>0$. In particular, given an $x\in\partial{\Omega}$, if there is a geodesic ray $\gamma$ such that $x\in\mathcal{C}(\gamma)$ then $x\in\overline{H^{b}_{o}(x,R)}$. Whether the later property holds true for a general complete Kobayashi hyperbolic domain $\Omega$ is not clear.
Lemma 3.3 also helps in understanding why for the bidisc ${\mathbb{D}}^{2}$, given $x=(x_{1},x_{2})\in\partial{{\mathbb{D}}^{2}}$, and $R>0$, $H^{b}_{o}(x,R)$, with pole 0, contains $x$ properly — as observed by Abate by direct computation. To see how, assuming $x_{1}\in\partial\mathbb{D}$, we consider a geodesic ray of ${\mathbb{D}}^{2}$ of the form $\gamma:=(\sigma,f\circ\sigma)$, where $\sigma$ is the radial geodesic ray of the unit disc starting at $0\in\mathbb{D}$ and landing at $x_{1}$ and $f$ is a holomorphic self-map of $\mathbb{D}$ such that the radial cluster set of $f$ at $x_{1}$ contains $x_{2}$ properly. It is a non-trivial fact that such an $f$ exist. Clearly then $\mathcal{C}(\gamma)$ contains $(x_{1},x_{2})$ properly and then by the above observation $H^{b}_{o}(x,R)$ contains $(x_{1},x_{2})$ properly.
For Kobayashi hyperbolic convex domains, we show that the closed big horospheres centered at a point contain that point. For this, we need a result due to Zimmer.
Result 3.4 (paraphrasing Zimmer [25, Lemma 3.2]).
Let $\Omega$ be a hyperbolic convex domain and let $o\in\Omega$. Given a boundary point $x\in\partial{\Omega}$, there exists $\alpha\geq 1$ and $\beta\geq 0$ such that the curve $\sigma_{x}:[0,\infty)\longrightarrow\Omega$ given by
$\sigma_{x}(t)=x+e^{-2t}(o-x)$
is an $(\alpha,\beta)$-quasi-geodesic.
We now present the result alluded to as above.
Proposition 3.5.
Let $\Omega$ be a Kobayashi hyperbolic convex domain and let $x\in\partial\Omega$. Then for any $R>0$ and any $o\in\Omega$, we have
$x\in\overline{H_{o}^{b}(x,R)}.$
Proof.
Suppose $x\notin\overline{{H_{o}^{b}(x,R)}}$ for some $R$ and $o\in\Omega$. Clearly, we can find an $r_{0}>0$ such that
$B(x,r_{0})\cap\Omega\cap\overline{H_{o}^{b}(x,R)}=\emptyset.$
Therefore, given $\epsilon>0$, for any $z\in B(x,r_{0})\cap\Omega$, we have
$\liminf_{w\to x}\big(\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(o,w)\big)\geq(1/2)\log R>(1/2)\log R-\epsilon.$
This implies that given $\epsilon>0$, and $z\in B(x,r_{0})\cap\Omega$, we can find a neighbourhood $U(z,\epsilon)$ of $x$ such that
$\inf_{w\in U(z,\epsilon)}\big(\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(o,w)\big)>({1}/{2})\log R-\epsilon$
which is equivalent to
$\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(o,w)>({1}/{2})\log R-\epsilon\ \ \forall w\in U(z,\epsilon).$ (3.2)
Let $\sigma_{x}:[0,\infty)\longrightarrow\Omega$ be an $(\alpha,\beta)$-quasi-geodesic as given by Result 3.4. In particular, if $z\in{\rm Range}({\sigma_{x}})\cap B(x,r_{0})$ and $w\in{\rm Range}({\sigma_{x}})\cap U(z,\epsilon)$, the above inequality is satisfied. Let $t_{2}>>t_{1}$ be such that $\sigma_{x}(t_{1})\in B(x,r_{0})$ and $\sigma_{x}(t_{2})\in U(z,\epsilon)$. Then
$\mathsf{k}_{\Omega}(\sigma_{x}(t_{1}),\sigma_{x}(t_{2}))\leq\alpha(t_{2}-t_{1})+\beta\ \ \text{and}\ \ \mathsf{k}_{\Omega}(o,\sigma_{x}(t_{2}))\geq\alpha t_{2}-\beta.$
This implies that
$\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(o,w)\leq 2\beta-\alpha t_{1}.$
Therefore, from (3.2), it follows that
$2\beta-\alpha t_{1}>({1}/{2})\log R-\epsilon\ \ \text{$\forall t_{1}$ such that $\sigma_{x}(t_{1})\in B(x,r_{0})$}.$
Since the set of all such $t_{1}$ is unbounded, we arrive at a contradiction, whence the result. ∎
We shall now present the proof of Proposition 1.2. In what follows, we shall use the following expression
$\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(o,w)=\mathsf{k}_{\Omega}(o,z)-2\langle z|w\rangle^{\Omega}_{o}.$
in the defining inequality of the horospheres. We now present:
The proof of Proposition 1.2.
Fix $o\in\Omega$ and $x\in\partial{\Omega}$. Since $\Omega$ is a visibility domain, there exists a geodesic ray $\gamma:[0,\infty)\longrightarrow\Omega$ that emanates from $o$ and lands at $x$, i.e.,
$\gamma(0)=o\ \ \ \text{and}\ \ \ \lim_{t\to\infty}\gamma(t)=x.$
The above result follows from [14, Lemma 2.16]; for bounded complete hyperbolic domains it was established in [10, Lemma 3.1]. Then by Lemma 3.3 for any $R>0$, we have $x\in\overline{H_{o}^{b}(x,R)}$.
Suppose, to get a contradiction, that there exists a $y\in\partial{\Omega}$ such that $y\neq x$ and $y\in\overline{H_{o}^{b}(x,R)}$ for some $R>0$. By definition, there exists a sequence $(y_{n})\subset H_{o}^{b}(x,R)$ such that $y_{n}\to y$ as $n\to\infty$. Note, this implies, for every $y_{n}$ we have
$\limsup_{w\to x}\big(2\langle y_{n}|w\rangle^{\Omega}_{o}-\mathsf{k}_{\Omega}(o,y_{n})\big)>({-1}/{2})\log R.$ (3.3)
Since $\Omega$ is a visibility domain, there exists a positive number $M$ such that
$\limsup_{z\to y,\,w\to x}\langle z|w\rangle^{\Omega}_{o}<M.$
The above inequality follows from [10, Proposition 2.4]. This implies there exists $n_{0}\in\mathbb{N}$ and an $r>0$ such that
$\sup_{w\in B(x,r)\cap\Omega}\langle y_{n}|w\rangle^{\Omega}_{o}<M\ \ \ \forall n\geq n_{0}.$
Therefore, for all $n\geq n_{0}$, we have
$\sup_{w\in B(x,r)\cap\Omega}\big(2\langle y_{n}|w\rangle^{\Omega}_{o}-\mathsf{k}_{\Omega}(o,y_{n})\big)<2M-\mathsf{k}_{\Omega}(o,y_{n}).$ (3.4)
From (3.3) and (3.4) we get that
$2M-\mathsf{k}_{\Omega}(o,y_{n})>({-1}/{2})\log R\Longleftrightarrow\mathsf{k}_{\Omega}(o,y_{n})<2M+({1}/{2})\log R\ \ \forall n\geq n_{0}.$
But this is a contradiction to the fact that $\lim_{n\to\infty}\mathsf{k}_{\Omega}(o,y_{n})=\infty$. Therefore, no point of $\partial\Omega$ other than $x$ is in the set $\overline{H_{o}^{b}(x,R)}\cap\partial{\Omega}$ whence the result. ∎
4. The proofs of Theorem 1.3, Theorem 1.5 Theorem 1.6
4.1. The proof of Theorem 1.3
Proof.
Fix a point $x\in\partial\Omega$ and let $o\in\Omega$ and $R<1$ be given. By [7, Theorem 1.11], we know that $(\Omega,\mathsf{k}_{\Omega})$ is Cauchy-complete and has visibility property with respect to the geodesics of the Kobayashi distance. Therefore, by Proposition 1.2, we have
$\overline{H_{o}^{s}(x,R)}\cap\partial\Omega\subset\overline{H_{o}^{b}(x,R)}\cap\partial{\Omega}=\{x\}.$
We shall now show that $x\in\overline{H_{o}^{s}(x,R)}\cap\partial\Omega$. Using visibility property, choose a geodesic ray $\beta:[0,\infty)\longrightarrow\Omega$ that starts from the point $o$ and lands at the point $x$, i.e. $\beta(0)=o$, $\lim_{t\to\infty}\beta(t)=x$. Choose $T$ large enough such that
$T+10\delta<(1/2)\log R,$ (4.1)
where $\delta\geq 0$ is such that $(\Omega,\mathsf{k}_{\Omega})$ is $\delta$-hyperbolic. Consider the point $z_{T}:=\beta(T)$.
Claim. For any $T$ that satisfies (4.1), $z_{T}\in H^{s}_{o}(x,R)$.
Fix a $T$ satisfying (4.1), and suppose, to get a contradiction, that the above claim is not true. Then we can find a sequence $(w_{n})\subset\Omega$ such that $w_{n}\to x$, as $n\to\infty$, and
$\lim_{n\to\infty}\big(\mathsf{k}_{\Omega}(z_{T},w_{n})-\mathsf{k}_{\Omega}(o,w_{n})\big)\geq(1/2)\log R.$ (4.2)
Let $\gamma_{n}:[0,L_{n}]\longrightarrow\Omega$ be a geodesic segment joining $o$ and $w_{n}$ and let $\sigma_{n}:[0,M_{n}]\longrightarrow\Omega$ be a geodesic segment joining $z_{T}$ and $w_{n}$. Appealing to Ascoli–Arzelá Theorem, and passing to subsequences, without loss of any generality, we can assume that $\gamma_{n}$, $\sigma_{n}$ converges to geodesic rays $\gamma$ and $\sigma$, respectively, locally uniformly on $0,\infty)$. Since $\Omega$ is a visibility domain, by [10, Lemma 3.1], both $\gamma$ and $\sigma$ land at the point $x$. By our assumption, $x$ represents a unique point in the Gromov boundary $\partial_{G}\Omega$, therefore both $\gamma$ and $\beta$ represent the same point in the Gromov boundary as represented by $x$. Since $\Omega$ is $\delta$-hyperbolic, using [12, Lemma 3.3, Chapter III.H], we have
$\mathsf{k}_{\Omega}(\gamma(t),\beta(t))\leq 2\delta\ \ \ \forall t\geq 0.$
Consider $\tilde{\beta}:[0,\infty)\longrightarrow\Omega$ defined by $\tilde{\beta}(t):=\beta(t+T)$ for all $t\geq 0$. Note that $\tilde{\beta}$ is a geodesic ray that starts at $z_{T}$ and lands at $x$. Therefore, by a similar argument as above, we get
$\mathsf{k}_{\Omega}(\sigma(t),\beta(t+T))\leq 2\delta\ \ \ \forall t\geq 0.$ (4.3)
From the above two inequalities we get:
$\mathsf{k}_{\Omega}(\sigma(t),\gamma(t+T))\leq\mathsf{k}_{\Omega}(\sigma(t),\beta(t+T))+\mathsf{k}_{\Omega}(\beta(t+T),\gamma(t+T))\leq 4\delta\ \ \ \forall t\geq 0.$ (4.4)
Fix a $t_{0}\geq 0$ and write
$\mathsf{k}_{\Omega}(z_{T},w_{n})=\mathsf{k}_{\Omega}(z_{T},\sigma_{n}(t_{0}))+\mathsf{k}_{\Omega}(\sigma_{n}(t_{0}),w_{n}).$
Now, for large $n$, we write
$\mathsf{k}_{\Omega}(o,w_{n})$ $=\mathsf{k}_{\Omega}(o,\gamma_{n}(t_{0}+T))+\mathsf{k}_{\Omega}(\gamma_{n}(t_{0}+T),w_{n})$
$\geq\mathsf{k}_{\Omega}(o,\sigma_{n}(t_{0}))-\mathsf{k}_{\Omega}(\sigma_{n}(t_{0}),\gamma_{n}(t_{0}+T))+\mathsf{k}_{\Omega}(w_{n},\sigma_{n}(t_{0}))-\mathsf{k}_{\Omega}(\sigma_{n}(t_{0}),\gamma_{n}(t_{0}+T))$
$=\mathsf{k}_{\Omega}(o,\sigma_{n}(t_{0}))-2\mathsf{k}_{\Omega}(\sigma_{n}(t_{0}),\gamma_{n}(t_{0}+T))+\mathsf{k}_{\Omega}(w_{n},\sigma_{n}(t_{0})).$
From the last two inequalities, for large enough $n$, we get
$\mathsf{k}_{\Omega}(z_{T},w_{n})-\mathsf{k}_{\Omega}(o,w_{n})\leq\mathsf{k}_{\Omega}(z_{T},\sigma_{n}(t_{0}))-\mathsf{k}_{\Omega}(o,\sigma_{n}(t_{0}))+2\mathsf{k}_{\Omega}(\sigma_{n}(t_{0}),\gamma_{n}(t_{0}+T)).$
Taking limit, as $n\to\infty$, and using inequality (4.4), we get
$\lim_{n\to\infty}\big(\mathsf{k}_{\Omega}(z_{T},w_{n})-\mathsf{k}_{\Omega}(o,w_{n})\big)$ $\leq\mathsf{k}_{\Omega}(z_{T},\sigma(t_{0}))-\mathsf{k}_{\Omega}(o,\sigma(t_{0}))+2\mathsf{k}_{\Omega}(\sigma(t_{0}),\gamma(t_{0}+T))$
$\leq t_{0}-\mathsf{k}_{\Omega}(o,\sigma(t_{0}))+8\delta$ (4.5)
Appealing to (4.3), we can write
$\mathsf{k}_{\Omega}(o,\sigma(t_{0}))$ $\geq\mathsf{k}_{\Omega}(o,\beta(t_{0}+T))-\mathsf{k}_{\Omega}(\beta(t_{0}+T),\sigma(t_{0}))$
$\geq t_{0}+T-2\delta$
Putting this into (4.1) and by our choice of $T$, we get
$\lim_{n\to\infty}\big(\mathsf{k}_{\Omega}(z_{T},w_{n})-\mathsf{k}_{\Omega}(o,w_{n})\big)$ $\leq t_{0}-(t_{0}+T-2\delta)+8\delta$
$=-T+10\delta$
$<(1/2)\log R,$
[…]
Therefore, given $x\in\partial\Omega$, $o\in\Omega$ and $R<1$, and a geodesic ray $\beta:[0,\infty)\longrightarrow\Omega$ that starts from $o$ and lands at $x$, for any $T>10\delta-(1/2)\log R$, the point $z_{T}=\beta(T)$ belongs to $H^{s}_{o}(x,R)$. In particular, $x\in\overline{H^{s}_{o}(x,R)}\cap\partial\Omega$. This establishes the result. ∎
We now present an alternative proof of Theorem 1.3 — as alluded to in Remark 1.4. However, we first recall a few definitions, and state two results that are at the heart of this proof.
Given $(\Omega,\mathsf{k}_{\Omega})$, a complete Kobayashi hyperbolic domain, that is Gromov hyperbolic, we shall denote by $\overline{\Omega}^{H}$, $\partial\overline{\Omega}^{H}$ the horofunction compactification, and the horofunction boundary respectively. We refer the reader to [4, Section 2] for a brief introduction in our setting to these objects. Given $\xi\in\partial\overline{\Omega}^{H}$, and $p\in\Omega$, by definition, there exists a unique continuous function $h_{p,\,\xi}$, vanishing at $p$, that represents $\xi$. Given $R>0$, the set
$H_{p}(\xi,R):=\left\{x\in X\,:\,h_{p,\,\xi}(x)<\tfrac{1}{2}\log R\right\}$
is called the horoball of radius $R$ centered at $\xi$ with pole $p$. It is known fact [24, Proposition 4.6] that there exists a unique $\Phi:\overline{\Omega}^{H}\longrightarrow\overline{\Omega}^{G}$ extending the identity map continuously on $\Omega$. Moreover, we have the following results:
Result 4.1 (paraphrasing [24, Proposition 4.4]).
Let $\xi_{1}$, $\xi_{2}\in\partial\overline{\Omega}^{H}$ be such that $\Phi(\xi_{1})=\Phi(\xi_{2})$. Then, for any $p\in\Omega$,
$\big|h_{p,\,\xi_{1}}(z)-h_{p,\,\xi_{2}}(z)\big|\leq 2\delta\quad\text{for all}\,z\in\Omega,$
where $\delta\geq 0$ is such that $(\Omega,\mathsf{k}_{\Omega})$ is $\delta$-hyperbolic.
Result 4.2 (paraphrasing [4, Proposition 6.5]).
Given $\xi\in\partial\overline{\Omega}^{H}$, $p\in\Omega$, for any $R>0$,
$\overline{H_{p}(\xi,R)}^{G}\cap\partial_{G}{\Omega}=\{\Phi(\xi)\}.$
Here, the closure at the left-hand side is in the Gromov topology.
Proof.
Fix a point $o\in\Omega$, $x\in\partial\Omega$ and let $R>0$ be given. Note $x$ represents a unique point in $\partial_{G}\Omega$. We claim that there exists an $R^{\prime}>0$ such that
$H_{o}(\xi,R^{\prime})\subset H^{b}_{o}(x,R^{\prime})\subset H^{s}_{o}(x,R)\subset H_{o}(\xi,R)\quad\forall\xi\in\Phi^{-1}(x).$
To see the claim, note that it is a fact that for any $S>0$
$H^{b}_{o}(x,S)=\bigcup_{\xi\in\Phi^{-1}(x)}H_{o}(\xi,S),\quad H^{s}_{o}(x,S)=\bigcap_{\xi\in\Phi^{-1}(x)}H_{o}(\xi,S).$
Therefore, we only need to show the second inclusion in the claim above. Let $R^{\prime}>0$ be such that $(1/2)\log{R^{\prime}}+2\delta\leq{(1/2)}\log{R}$. Let $z\in H^{b}_{o}(x,R^{\prime})$, then there exists $\xi\in\Phi^{-1}(x)$ such that $z\in H_{o}(\xi,R^{\prime})$. Now if $\zeta\in\Phi^{-1}(x)$ be any other point, then using Result 4.1 we get
$h_{o,\,\zeta}(z)\leq 2\delta+h_{o,\,\xi}(z)<2\delta+(1/2)\log R^{\prime}\leq(1/2)\log R.$
Therefore, for any $\zeta\in\Phi^{-1}(x)$, we have $z\in H_{p}(\zeta,R)$ whence $H^{b}_{o}(x,R^{\prime})\subset H^{s}_{o}(x,R)$. This establishes the claim.
The claim above together with Result 4.2 implies that $\overline{H^{s}_{o}(x,R)}^{G}\cap\partial_{G}\Omega=\{x\}$. For big horospheres, one could change the role of $R,R^{\prime}$ in the claim above, and then the similar argument works. ∎
4.2. The proof of Theorem 1.5
We shall now present the proof of Theorem 1.5, but first we recall: given $\Omega$, a complete Kobayashi hyperbolic domain, we say that $\Omega$ is a local model-Gromov-hyperbolic domain if for any point $x\in\partial\Omega$ there is a neighbourhood $U_{x}$ in $\mathbb{C}^{d}$ such that $\Omega\cap U_{x}$ is connected and $(\Omega\cap U_{x},\mathsf{k}_{\Omega\cap U_{x}})$ is a model-Gromov-hyperbolic domain.
The proof of Theorem 1.5.
Fix an $x\in\partial{\Omega}$, then there exists a neighbourhood $U_{x}$ of $x$ in $\mathbb{C}^{d}$ such that $(\Omega\cap U_{x},\mathsf{k}_{\Omega\cap U_{x}})$ is a model-Gromov-hyperbolic domain. Denote by $V_{x}:=U_{x}\cap\Omega$. By [7, Theorem 1.11], it follows that $V_{x}$ has the weak visibility property with respect to $\mathsf{k}_{V_{x}}$. Note that since $x\in\partial\Omega$ is arbitrary, and $\Omega$ is hyperbolically embedded in $\mathbb{C}^{d}$, appealing to Result 2.4, we get that $\Omega$ is a weak visibility domain. Since $\Omega$ is complete Kobayashi hyperbolic, by Result 2.3, $\Omega$ also has visibility property with respect to the geodesics of the Kobayashi distance. Therefore, by Proposition 1.2, we have
$\overline{H_{o}^{s}(x,R)}\cap\partial\Omega\subset\overline{H_{o}^{b}(x,R)}\cap\partial{\Omega}=\{x\}.$
We shall now show that $x\in\overline{H_{o}^{s}(x,R)}\cap\partial\Omega$ for all $R>0$. Given $x\in\partial\Omega$, we choose $r$ small enough such that $B(x,r)\Subset U_{x}$. Let $W_{x}:=B(x,r)\cap\Omega$. Since $\Omega$ is a visibility domain, by Result 2.5, there exists a $C>0$ such that
$\mathsf{k}_{\Omega}(z,w)\leq\mathsf{k}_{V_{x}}(z,w)\leq\mathsf{k}_{\Omega}(z,w)+C\ \ \ \forall z,w\in W_{x}.$
Fix a point $p\in W_{x}$, then for a point $z\in W_{x}$, we have
$\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(p,w)\leq\mathsf{k}_{V_{x}}(z,w)-\mathsf{k}_{V_{x}}(p,w)+C,\ \ \ \forall w\in W_{x}.$
The above inequality then implies that for $z,p\in W_{x}$ we have
$\limsup_{w\to x}\big(\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(p,w)\big)\leq\limsup_{w\to x}\big(\mathsf{k}_{V_{x}}(z,w)-\mathsf{k}_{V_{x}}(p,w)\big)+C.$
Therefore, it follows that
$W_{x}\cap H^{s}_{p}(x,R,{V_{x}})\subset H^{s}_{p}(x,Re^{2C},\Omega),$ (4.6)
where $p\in W_{x}$ is fixed. Now, since $({V_{x}},\mathsf{k}_{V_{x}})$ is a model-Gromov-hyperbolic domain by Theorem 1.3 we have
$x\in\overline{H^{s}_{p}(x,R,{V_{x}})}\ \ \ \forall R>0.$
Therefore, it follows from (4.6) that for all $R>0$, we have
$x\in\overline{H^{s}_{p}(x,Re^{2C},\Omega)}.$
Since $R$ is arbitrary, it follows that $x\in H^{s}_{p}(x,R)$ for all $R>0$ and then arguing as at the end of the proof of Lemma 3.3, we have $x\in H^{s}_{o}(x,R)$ for all $R>0$, $o\in\Omega$ and $x\in\partial{\Omega}$. This establishes the result. ∎
4.3. The proof of Theorem 1.6
In this subsection, we shall present the proof of Theorem 1.6 but first we need a lemma.
Lemma 4.3.
Let $\Omega$ be a bounded convex domain and let $x\in\partial\Omega$ be a Dini-smooth boundary point. Then for any $R>0$ and $o\in\Omega$ we have
$x\in\overline{H^{s}_{o}(x,R)}.$
Proof.
Since $\Omega$ is a bounded convex domain, by [20, Proposition 2.4], there exists a $C\in\mathbb{R}$ such that for every $w\in\Omega$, we have
$\mathsf{k}_{\Omega}(o,w)\geq C+\frac{1}{2}\log\frac{1}{\delta_{\Omega}(w)},$
where $\delta_{\Omega}(\cdot)$ denotes the boundary distance function. Since $x\in\partial{\Omega}$ is a Dini-smooth boundary point, by a result of Nikolov–Andreev [21], there exists $\epsilon>0$, $c>0$, such that for every $z,w\in B(x,\epsilon)$, we have
$\mathsf{k}_{\Omega}(z,w)\leq\log\Bigg(1+\frac{c\|z-w\|}{\sqrt{\delta_{\Omega}(z)\delta_{\Omega}(w)}}\Bigg).$
From the above two inequalities, for any $z,w\in B(x,\epsilon)$ we get
$\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(o,w)\leq\log\Bigg(\sqrt{\delta_{\Omega}(w)}+\frac{c\|z-w\|}{\sqrt{\delta_{\Omega}(z)}}\Bigg)-C.$
Therefore, for a fixed $z\in B(x,\epsilon)$, we have
$\limsup_{w\to x}\big(\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(o,w)\big)$ $\leq\log\Bigg(\frac{c\|z-x\|}{\sqrt{\delta_{\Omega}(z)}}\Bigg)-C$
$=\frac{1}{2}\log{\|z-x\|}+\frac{1}{2}\log\frac{\|z-x\|}{\delta_{\Omega}(z)}+(\log c-C).$
If we choose $z$ on the normal line at the point $x$, and $z$ is sufficiently close to $x$, then the second term in the inequality above remains bounded above by a constant. This shows for any $R<1$, we can find an $s>0$ such that for any point $z$ that lies on the normal line passing through $x$ with $\delta_{\Omega}(z)<s$, we have
$\limsup_{w\to x}\big(\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(o,w)\big)<\frac{1}{2}\log R.$
From this it follows that for any $R<1$, and hence for any $R>0$, the set $\overline{H^{s}_{o}(x,R)}\cap\partial{\Omega}$ contains $x$. This together with our observation at the start establishes the result. ∎
We now present
The proof of 1.6.
Since $\Phi_{x}(V_{x}\cap\Omega)$ has visibility property with respect to the geodesics and $\Phi_{x}$ is a biholomorphism, it is not difficult to check that $V_{x}\cap\Omega$ also has the visibility property with respect to the geodesics of $\mathsf{k}_{V_{x}\cap\Omega}$. By Result 2.3, $(V_{x}\cap\Omega,\mathsf{k}_{V_{x}\cap\Omega})$ is a weak visibility domain. Since every bounded domain is hyperbolically imbedded in $\mathbb{C}^{d}$, appealing to Result 2.4, we deduce that $(\Omega,\mathsf{k}_{\Omega})$ is a weak visibility domain. Since $\mathsf{k}_{\Omega}$ is complete, it satisfies visibility property with respect to the geodesics. Therefore, by Proposition 1.2, given $o\in\Omega$, $x\in\partial\Omega$ and $R>0$ we have
$\overline{H^{s}_{o}(x,R)}\cap\partial{\Omega}\subset\overline{H^{b}_{o}(x,R)}\cap\partial{\Omega}=\{x\}.$
Fix a point $x\in\partial\Omega$, to show that $x\in\overline{H^{s}_{o}(x,R)}$, consider $(\Phi_{x},U_{x},V_{x})$ as in the definition of local convexity. Then, for very small $r$, by the localization result, there exists a $C>0$ such that
$\mathsf{k}_{\Omega}(z,w)\leq\mathsf{k}_{V_{x}\cap\Omega}(z,w)\leq\mathsf{k}_{\Omega}(z,w)+C\ \ \ \forall\ z,w\in B(x,r)\cap\Omega.$
Write $W_{x}:=B(x,r)\cap\Omega$. Fix a point $p\in W_{x}$, then for a point $z\in W_{x}$, we have
$\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(p,w)\leq\mathsf{k}_{V_{x}\cap\Omega}(z,w)-\mathsf{k}_{V_{x}\cap\Omega}(p,w)+C,\ \ \ \forall w\in W_{x}.$
The above inequality then implies that for $z,p\in W_{x}$ we have
$\limsup_{w\to x}\big(\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(p,w)\big)\leq\limsup_{w\to x}\big(\mathsf{k}_{V_{x}\cap\Omega}(z,w)-\mathsf{k}_{V_{x}\cap\Omega}(p,w)\big)+C.$
Therefore it follows that
$W_{x}\cap H^{s}_{p}(x,R,{V_{x}}\cap\Omega)\subset H^{s}_{p}(x,Re^{2C},\Omega),$ (4.7)
where $p\in W_{x}$ is fixed. Note that $\Phi_{x}(V_{x}\cap\Omega)$ is a bounded convex domain and $\Phi_{x}(x)$ is a Dini-smooth boundary point. Therefore, by Lemma 4.3,
$\Phi_{x}(x)\in\overline{H^{s}_{\Phi_{x}(p)}(\Phi_{x}(x),R,\Phi_{x}({V_{x}}\cap\Omega))}.$
Since $\Phi_{x}:V_{x}\cap\Omega\longrightarrow\Phi_{x}({V_{x}}\cap\Omega)$ is a biholomorphism that extends as a homeomorphism, we get that
$x\in\overline{H^{s}_{p}(x,R,{V_{x}}\cap\Omega)}\ \ \ \forall R>0.$
Therefore, it follows from (4.7) that for all $R>0$, we have
$x\in\overline{H^{s}_{p}(x,Re^{2C},\Omega)}.$
Since $R$ is arbitrary, it follows that $x\in H^{s}_{p}(x,R,\Omega)$ for all $R>0$ and then arguing as at the end of the proof of Lemma 3.3, we have $x\in H^{s}_{o}(x,R,\Omega)$ for all $R>0$, $o\in\Omega$ and $x\in\partial{\Omega}$. This establishes the result. ∎
5. Continuous extension of biholomorphisms
In this section, we present the proof of our result regarding the continuous extension of biholomorphisms. As an upshot, we shall characterize those bounded simply-connected domains in $\mathbb{C}$ that have property (1.1) in problem ()(*). We first recall a Julia lemma essentially due to Abate.
Result 5.1.
Let $D_{1}\subset\mathbb{C}^{n}$ and $D_{2}\subset\mathbb{C}^{m}$ be bounded domains, and let $f:D_{1}\to D_{2}$ be a holomorphic map. Suppose there exist points $p\in D_{1}$, $q\in D_{2}$, a constant $\alpha>0$, and a sequence $(w_{\nu})\subset D_{1}$ converging to $x\in\partial D_{1}$ such that $f(w_{\nu})\to y\in\partial D_{2}$ and
$\lim_{\nu\to\infty}\!\big(\mathsf{k}_{D_{1}}(p,w_{\nu})-\mathsf{k}_{D_{2}}(q,f(w_{\nu}))\big)\leq\tfrac{1}{2}\log\alpha.$
Then, for every $R>0$,
$f\big(H^{s}_{p}(x,R)\big)\subset H^{b}_{q}(y,\alpha R).$
Abate proved the above result when $D_{1}=D_{2}$ and $p=q$ (see, e.g., [2, Proposition 2.4.15]). The general case follows by the same argument.
Lemma 5.2.
Let $D$, $\Omega$ be bounded domains that are complete with respect to the Kobayashi distance. Suppose there exist $o\in D$, $p\in\Omega$ such that
  • for any two points $y_{1}\neq y_{2}\in\partial D$, there exists an $R>0$ such that $H^{b}_{o}(y_{1},R)\cap H^{b}_{o}(y_{2},R)=\emptyset$, and
  • for any $x\in\partial\Omega$ and any $R>0$, $H^{s}_{p}(x,R)\neq\emptyset$.
Then any biholomorphism $\Phi:\Omega\to D$ extends as a continuous map from $\overline{\Omega}$ onto $\overline{D}$.
Proof.
Let $o\in D$ be as in the statement of the lemma. Observe that — owing to the last part of Result 2.1 — we may assume that $p\in\Omega$ satisfies $\Phi(p)=o$. We claim that, given $x\in\partial\Omega$, let $(w_{n})\subset\Omega$ be a sequence that converges to $x$ such that $\Phi(w_{n})\to y\in\partial D$. Then
$\Phi(H^{s}_{p}(x,R))\subset H^{b}_{o}(y,R)\ \ \ \forall R>0.$ (5.1)
To see this, we note that — since $\Phi$ is an isometry — for each $n$, we have
$\mathsf{k}_{\Omega}(p,w_{n})-\mathsf{k}_{D}(o,\Phi(w_{n}))=\mathsf{k}_{\Omega}(p,w_{n})-\mathsf{k}_{D}(\Phi(p),\Phi(w_{n}))=\mathsf{k}_{\Omega}(p,w_{n})-\mathsf{k}_{D}(p,w_{n})=0.$
Therefore, an application of Result 5.1 with $\alpha=1$, establishes (5.1).
Fix a point $x\in\partial\Omega$, we shall show that the limit
$\lim_{z\in\Omega,\,z\to x}\Phi(z)$
exists. Suppose, to get a contradiction, that the above limit does not exist. This implies that — since $\Phi$ is a biholomorphism and $D$ is bounded — there exist sequences $(z_{n})$ and $(w_{n})$ in $\Omega$ that converge to $x$ and points $y_{1}\neq y_{2}\in\partial D$ such that $\Phi(z_{n})\to y_{1}$, $\Phi(w_{n})\to y_{2}$. Applying (5.1), we get
$\Phi(H^{s}_{o}(x,R))\subset H^{b}_{o}(y_{1},R)\cap H^{b}_{o}(y_{2},R)\ \ \ \forall R>0.$
Since $y_{1}$ and $y_{2}$ are distinct, there exists an $R_{0}>0$ such that $H^{b}_{o}(y_{1},R)\cap H^{b}_{o}(y_{2},R)=\emptyset$ for all $R<R_{0}$. This implies that $H^{s}_{o}(x,R)=\emptyset$ for all $R<R_{0}$ which is a contradiction to our hypothesis whence the result.
It is now routine to check that $\Phi$ extends continuously on $\overline{\Omega}$. ∎
We now present
The proof of Theorem 1.9.
Since $D$ is metrically-regular, there exists a point $o\in D$ such that $(1)$ and $(2)$ in Definition 1.7 are satisfied. Let $\Phi:\Omega\longrightarrow D$ be the inverse of the map $\Psi$ and let $p\in\Omega$ be such that $\Phi(p)=o$.
Lemma 5.2 implies that $\Phi:\Omega\longrightarrow D$ extends as a continuous map from $\overline{\Omega}$ onto $\overline{D}$ — that we continue to denote as $\Phi$ itself. Consider a point $x\in\partial\Omega$ and let $\xi=\Phi(x)$. Given a point $z\in H^{b}_{p}(x,R)$, there exists a sequence $(z_{n})$ in $\Omega$ such that $z_{n}\to x$ and
$\liminf_{w\to x}\big(\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(p,w)\big)=\lim_{n\to\infty}\big(\mathsf{k}_{\Omega}(z,z_{n})-\mathsf{k}_{\Omega}(p,z_{n})\big)<(1/2)\log R.$
Since $\Phi$ is an isometry with respect to the Kobayashi distances, from the above inequality, we get
$\lim_{n\to\infty}\big(\mathsf{k}_{D}(\Phi(z),\Phi(z_{n}))-\mathsf{k}_{D}(o,\Phi(z_{n}))\big)<(1/2)\log R.$
Note $\Phi$ is continuous at $x$ and $\Phi(x)=\xi$, since $(z_{n})$ converges to $x$, it follows that $(\Phi(z_{n}))$ converges to $\xi$. This implies that $\Phi(z)\in H(\xi,R)$ for any $z\in H^{b}_{p}(x,R)$. Therefore,
$\Phi(H^{b}_{p}(x,R))\subset H(\Phi(x),R)\ \ \ \forall R>0.$ (5.2)
Since small horospheres and big horospheres in $D$ are identical, and $\Psi$ an isometry, an application of Result 5.1 — as in the proof of Lemma 5.2 — implies that given $\zeta\in\partial D$, and $(\zeta_{n})\subset D$, a sequence converging to $\zeta$, such that $\Psi(\zeta_{n})\to y$ for some $y\in\partial\Omega$, we have
$\Psi(H(\zeta,R)\subset H^{b}_{p}(y,R)\ \ \ \forall R>0.$ (5.3)
Fix a point $\xi\in\partial D$, we shall show that the limit
$\lim_{z\in D,\,z\to\xi}\Psi(z)\ \ \text{exists}.$
Suppose, to get a contradiction, that the above is not true. This implies that — since $\Psi$ is a biholomorphism and $\Omega$ is bounded — there exist sequences $(z_{n}),(w_{n})\subset D$ converging to $\xi$ and $x_{1}\neq x_{2}\in\partial\Omega$ such that $\Psi(z_{n})\to x_{1}$ and $\Psi(w_{n})\to x_{2}$. Applying 5.3, we get
$\Psi(H(\xi,R))\subset H^{b}_{p}(x_{i},R)\ \ \ \forall R>0,\ i=1,2.$
Since $\Phi$ is continuous at $x_{1}$, the sequence $\Phi(\Psi(z_{n}))=z_{n}$ converges to $\Phi(x_{1})$, as $n\to\infty$. This implies that $\Phi(x_{1})=\xi$. Similarly $\Phi(x_{2})=\xi$. By (5.2), we have
$\Phi(H^{b}_{p}(x_{i},R))\subset H(\xi,R)\ \ \ \forall R>0,\,i=1,2.$
From the last two relations, for each $i=1,2$, we get
$H^{b}_{p}(x_{i},R)\subset\Psi(H(\xi,R))\subset H^{b}_{p}(x_{i},R)\ \ \ \forall R>0.$
This implies that $H^{b}_{p}(x_{1},R)=H^{b}_{p}(x_{2},R)=\Psi(H(\xi,R))$ for all $R>0$. But this contradicts our hypothesis regarding the big horospheres on the domain $\Omega$.
This implies that $\Psi$ extends as a continuous map from $\overline{D}$ onto $\overline{\Omega}$. Therefore $\Psi:D\longrightarrow\Omega$ extends as a homeomorphism from $\overline{D}$ onto $\overline{\Omega}$. ∎
As a corollary of the above result, we present:
The Proof of Theroem 1.1.
Recall $\Omega$ is a bounded simply-connected domain that satisfies
$\overline{H^{s}_{o}(x,R)}\cap\partial\Omega=\overline{H^{b}_{o}(x,R)}\cap\partial\Omega=\{x\},$
for any $x\in\partial\Omega$ and for any $R>0$, with $o=\psi(0)$, where $\psi:\mathbb{D}\longrightarrow\Omega$ is a Riemann map. Clearly $H^{s}_{o}(x,R)$ is nonempty for all $x\in\partial\Omega$, $R>0$ and also given $x\neq y\in\partial\Omega$, $H^{b}_{o}(x,R)\neq H^{b}_{o}(y,R)$ for any $R>0$. Clearly $\mathbb{D}$ is metrically-regular, therefore, by Theorem 1.9 — taking $D=\mathbb{D}$ — we see that $\psi$ extends as a homeomorphism from $\overline{\mathbb{D}}$ onto $\overline{\Omega}$. Hence $\Omega$ has to be a Jordan domain. ∎
We now present an example of a domain that does not satisfy condition $(a)$ in Theorem 1.9 but satisfies condition $(b)$ therein.
Example 5.3.
Consider the simply-connected domain
$\Omega:=\mathbb{D}\setminus\{[0,1)\}.$
Let $\psi:\mathbb{D}\longrightarrow\Omega$ be a Riemann map and $\phi:\Omega\longrightarrow\mathbb{D}$ be its inverse. Since $\partial\Omega$ is locally connected, by Carathéodory’s Extension Theorem [8, Theorem 4.3.1], $\psi$ has a continuous extension $\hat{\psi}$ from $\overline{\mathbb{D}}$ onto $\overline{\Omega}$. Then, for a given $x\in(0,1]$, $\hat{\psi}^{-1}(x)$ has two pre-images, say $\xi_{1}\neq\xi_{2}\in\partial\mathbb{D}$, see [8, Proposition 4.3.5]. Proceeding exactly as in the claim that appears in the proof of Lemma 5.2, we get that
$\phi\big(H^{s}_{o}(x,R)\big)\subset H(\xi_{1},R)\cap H(\xi_{2},R)\ \ \ \forall R>0,$
where $o=\psi(0)$. As $\xi_{1}$ and $\xi_{2}$ are distinct points in $\partial\mathbb{D}$, we can find an $R_{0}>0$ such that for any $R<R_{0}$, $H(\xi_{1},R)\cap H(\xi_{2},R)=\emptyset$. Therefore, for each $x\in(0,1]$, there exists an $R_{0}>0$ such that $H^{s}_{o}(x,R)=\emptyset$ for all $R<R_{0}$. Consequently, $\Omega$ does not satisfy condition $(a)$ in Theorem 1.9.
It is a fact that simply-connected visibility domains are exactly those whose boundary is locally connected, therefore, $\Omega$ is a visibility domain and hence satisfies
$\overline{H^{b}_{o}(x,R)}\cap\partial\Omega=\{x\}\ \ \ \forall R>0.$
This clearly implies that $\Omega$ satisfies condition $(b)$ in Theorem 1.9.
Remark 5.4.
Let $\Omega$ be a bounded simply-connected domain and let $\phi:\Omega\longrightarrow\mathbb{D}$ be a biholomorphism. Suppose $\Omega$ has the property that, given $o\in\Omega$, for any $x\in\partial\Omega$, and $R>0$, $H^{s}_{o}(x,R)$ is nonempty. Then, by Lemma 5.2, $\phi$ extends continuously from $\overline{\Omega}$ onto $\overline{\mathbb{D}}$. Then it is a fact that $\phi$ has to be a homeomorphism [15, Theorem 5.12] and hence $\Omega$ is a Jordan domain. Therefore, for a bounded simply-connected domain $\Omega$ we have the following dichotomy: either they are Jordan domains or there exists a point $x\in\partial\Omega$ and an $R>0$ such that $H^{s}_{o}(x,R)=\emptyset$ for some $o\in\Omega$.
6. Examples
In one dimension, we introduce a class of domains that are locally a model-Gromov-hyperbolic, in particular, they also include non-Gromov-hyperbolic domains and domains whose boundary could be a fractal set.
Definition 6.1.
Let $\Omega\subset\mathbb{C}$ be a Kobayashi hyperbolic domain. We say that $\Omega$ satisfies Condition 1 if for any point $p\in\partial{\Omega}$, there exist an $r>0$ and a topological embedding $\tau_{p}:\overline{\mathbb{D}}\longrightarrow\overline{\Omega}$ such that $\tau_{p}(\mathbb{D})\subset\Omega$ and $B(p;r)\cap\Omega\subset\tau_{p}(\mathbb{D})$.
It is easy to see why domains that satisfy Condition 1 are local-model-Gromov-hyperbolic domain: take a point $p\in\partial\Omega$, and let $\tau_{p}$, $r>0$ be as above in the definition. Note that $\tau_{p}(\mathbb{D})$ is a Jordan domain, therefore appealing to the Riemann Mapping Theorem and Carathéodory’s Extension Theorem, we can assume, without loss of any generality, that $\tau_{p}$ is a biholomorphism that extends as a homeomorphism from $\mathbb{D}$ to $\overline{\tau_{p}(\mathbb{D})}$. Clearly $\tau_{p}(\mathbb{D})$ equipped with its Kobayashi distance is a model-Gromov-hyperbolic metric space. Taking
$U_{p}:=\tau_{p}(\mathbb{D})\cup B(p,r)$
we see that $U_{p}\cap\Omega=\tau_{p}(\mathbb{D})$, since $p\in\Omega$ is arbitrary, we see that $\Omega$ is a local model-Gromov-hyperbolic domain.
Planar domains that satisfy Condition 1 were recently introduced in [14] where a characterization of such domains is also given: these are domains for which boundary is locally connected and each boundary component is a Jordan curve in the Riemann sphere [14, Proposition 4.15]. It was also established that these domains satisfy the strongest form of visibility property.
Example 6.2 ([7, Proposition 9.1]).
Consider the domain $\Omega\subset\mathbb{C}$ defined by
$\Omega:=\mathbb{C}\setminus\cup_{m,n\in\mathbb{Z}}\{z\in\mathbb{C}\,:\,|z-(m+ni)|<1/4\}.$
Clearly $\Omega$ satisfies Condition 1 but $(\Omega,\mathsf{k}_{\Omega})$ is not Gromov hyperbolic. The obstruction to Gromov hyperbolicity comes from the existence of a $\mathbb{Z}^{2}$ action, see [7, Proposition 9.1] for the details.
The second example is an example of a domain where each of its boundary component is a Jordan curve given locally by the graph of a nowhere differentiable continuous function. It is borrowed from [14, Example 6.6].
Example 6.3.
The Takagi or blancmange function is a continuous nowhere-differentiable function defined as follows.
$\forall\,t\in\mathbb{R},\;T(t):=\sum_{j=0}^{\infty}2^{-j}\mathrm{dist}(2^{j}t,\mathbb{Z}).$ (6.1)
[…]
Acknowledgments
The authors are grateful to the anonymous referee for several helpful comments and suggestions, particularly for pointing out an alternative proof of Theorem 1.3 and for remarks that improved the clarity of Section 5.